y=2x-8 plot the graphs of the given linear equations using the x and y intercepts and a checkpoint. I don't get it??
Answer by nyc_function(2741)

Let x = 0 and solve for y. This will give you one point.

Then let y = 0 and solve for x, this will give you the second point.
=====================================================================
Let x = 0

y = 2x - 8

y = 2(0) - 8

y = 0 - 8

y = -8

The first point to plot is (0,-8).


===================================

Now let y = 0 and solve for x.

y = 2x - 8

0 = 2x - 8

8 = 2x

8/2 = x

4 = x

The second point to plot is (4,0).

Plot both points on graph paper and then connect the points with a straight line.

Copy and paste the link below to see your graph.

http://mathway.com/answer.aspx?p=grap?p=ySMB012x-8?p=False?p=False?p=True?p=False?p=True